# Depth Priors in Removal Neural Radiance Fields

### v2

@@ 4.3.1 Depth Evaluation on KITTI Datasets
This set of experiments is meant to support our first assumption that the depth generated by COLMAP can be taken as depth ground truth in scenarios lacking ground truth information. This will further support our second assumption that we can take depth generated by COLMAP as ground truth to evaluate depth obtained by monocular depth estimation models.
We conducted rigorous evaluations of depth maps from COLMAP’s dense depth reconstruction and various monocular depth estimation methods using the selected KITTI subsets with ground truth depth maps. The results, presented in Table 1, reveal that COLMAP consistently outperforms monocular methods across all metrics, with a $\delta_{1}$ value of 91%, highlighting its closer approximation to ground truth. In scenarios where acquiring ground truth is cost-prohibitive, COLMAP’s dense depth reconstruction proves to be a viable alternative. Additional details and visual comparisons of depth maps generated by different methods are illustrated in Figure 2. For instance, the COLMAP dense depth map, as shown in the second row of Figure 2, reveals a more precise delineation and representation of object edges compared to maps produced by monocular depth estimation methods.
@@
Methods RMSE $\downarrow$ $\delta_{1}\uparrow$ $\log_{10}\downarrow$
COLMAP 0.017 0.910 0.042
EcoDepth 0.017 0.670 0.080
ZoeDepth 0.024 0.807 0.064
Depth Anything 0.044 0.184 0.357
@@ 4.3.2 Depth Evaluation on Removal NeRF Dataset
ZoeDepth RMSE $\downarrow$ 0.102 0.034 0.026 0.074 0.035 0.045 0.032 0.064 0.035 0.063
$\delta_{1}\uparrow$ 0.526 0.994 0.997 0.946 0.984 0.999 0.993 0.989 0.989 0.989
$\log_{10}\downarrow$ 0.146 0.026 0.024 0.290 0.016 0.020 0.021 0.028 0.017 0.023
Density 0.968 0.991 0.991 0.887 0.981 0.987 0.983 0.849 0.911 0.811
Depth Anything RMSE $\downarrow$ 0.173 0.305 0.287 0.264 0.449 0.325 0.325 0.638 0.356 0.496
$\delta_{1}\uparrow$ 0.058 0.257 0.225 0.325 0.363 0.274 0.247 0.154 0.218 0.261
$\log_{10}\downarrow$ 0.636 0.213 0.207 0.196 0.189 0.159 0.250 0.341 0.222 0.216
Density 0.968 0.991 0.991 0.887 0.981 0.987 0.983 0.849 0.911 0.811
ECoDepth RMSE $\downarrow$ 0.075 0.054 0.032 0.093 0.065 0.073 0.106 0.122 0.062 0.168
$\delta_{1}\uparrow$ 0.496 0.953 0.994 0.824 0.962 0.920 0.621 0.893 0.939 0.625
$\log_{10}\downarrow$ 0.130 0.045 0.022 0.058 0.037 0.032 0.084 0.048 0.036 0.083
Density 0.968 0.991 0.991 0.887 0.981 0.987 0.983 0.849 0.911 0.811

### v3

@@ 4.3.1 Depth Evaluation on KITTI Datasets
This set of experiments is meant to support our first assumption that the depth generated by COLMAP can be taken as depth ground truth in scenarios lacking ground truth information. This will further support our second assumption that we can take depth generated by COLMAP as ground truth to evaluate depth obtained by monocular depth estimation models. All experiments were conducted using an NVIDIA A100 GPU.
We conducted rigorous evaluations of depth maps from COLMAP’s dense depth reconstruction and various monocular depth estimation methods using the selected KITTI subsets with ground truth depth maps. The results, presented in Table 1, reveal that COLMAP consistently outperforms (or achieves on-the-par results) monocular methods across all metrics, with a $\delta_{1}$ value of 91%, highlighting its closer approximation to ground truth. In scenarios where acquiring ground truth is cost-prohibitive, COLMAP’s dense depth reconstruction proves to be a viable alternative. Additional details and visual comparisons of depth maps generated by different methods are illustrated in Figure 2. For instance, the COLMAP dense depth map, as shown in the second row of Figure 2, reveals a more precise delineation and representation of object edges compared to maps produced by monocular depth estimation methods.
@@
Methods ${\delta_{1}\uparrow}$ RMSE $\downarrow$ $\log_{10}\downarrow$
COLMAP 0.910 0.017 0.042
ZoeDepth 0.807 0.024 0.064
EcoDepth 0.670 0.017 0.080
Depth Anything 0.184 0.044 0.357
@@ 4.3.2 Depth Evaluation on Removal NeRF Dataset
ZoeDepth ${\delta_{1}\uparrow}$ 0.53 0.99 1.00 0.95 0.98 1.00 0.99 0.99 0.99 0.99
RMSE $\downarrow$ 0.10 0.03 0.03 0.07 0.04 0.05 0.03 0.06 0.04 0.06
$\log_{10}\downarrow$ 0.15 0.03 0.02 0.29 0.02 0.02 0.02 0.03 0.02 0.02
Density 0.97 0.99 0.99 0.89 0.98 0.99 0.98 0.85 0.91 0.81
Depth Anything $\delta_{1}\uparrow$ 0.06 0.26 0.23 0.33 0.36 0.27 0.25 0.15 0.22 0.26
RMSE $\downarrow$ 0.17 0.31 0.29 0.26 0.45 0.33 0.33 0.64 0.36 0.50
$\log_{10}\downarrow$ 0.64 0.21 0.21 0.20 0.19 0.16 0.25 0.34 0.22 0.22
Density 0.97 0.99 0.99 0.89 0.98 0.99 0.98 0.85 0.91 0.81
ECoDepth $\delta_{1}\uparrow$ 0.50 0.95 0.99 0.82 0.96 0.92 0.62 0.89 0.94 0.63
RMSE $\downarrow$ 0.08 0.05 0.03 0.09 0.07 0.07 0.11 0.12 0.06 0.17
$\log_{10}\downarrow$ 0.13 0.05 0.02 0.06 0.04 0.03 0.08 0.05 0.04 0.08
Density 0.97 0.99 0.99 0.89 0.98 0.99 0.98 0.85 0.91 0.81